# Evaluation for TL;DR

### v2

@@ Abstract
Many decision-making tasks, where both accuracy and efficiency matter, still require human supervision. For example, tasks like traffic officers reviewing hour-long dashcam footage or researchers screening conference videos can benefit from concise summaries that reduce cognitive load and save time. Yet current vision-language models (VLMs) often produce verbose, redundant outputs that hinder task performance. Existing video caption evaluation depends on costly human annotations and overlooks the summaries’ utility in downstream tasks. We address these gaps with Video-to-text Information Bottleneck Evaluation (VIBE), an annotation-free method that scores VLM outputs using two metrics: grounding (how well the summary aligns with visual content) and utility (how informative it is for the task). VIBE selects from randomly sampled VLM outputs by ranking them according to the two scores to support effective human decision-making. Human studies on LearningPaper24, SUTD-TrafficQA, and LongVideoBench show that summaries selected by VIBE consistently improve performance—boosting task accuracy by up to $61.23\%$ and reducing response time by $75.77\%$ compared to naive VLM summaries or raw video. [^1] Source code: https://github.com/UTAustin-SwarmLab/Task-aware-TLDR-Public
@@ 4 Video-to-Text Information Bottleneck Evaluation (VIBE)
We condition both the numerator and denominator with $T_{\mathrm{masked}}$, a masked version of the text, to approximate $\mathbf{I}^{P}(V;T)$. The approximation holds if $T_{\mathrm{masked}}$ is independent of both the video $V$ and the original text $T$. Masking key content makes $T_{\mathrm{masked}}$ effectively uninformative and thus independent. Following [20], we use tf-idf [37] to identify keywords and replace them with a "<MASK>" token. This masking helps ensure that eq.˜3 reflects the true dependency between $V$ and $T$, rather than exploiting shortcuts from the text itself. Thus, the grounding score measures how much the TL;DR remains genuinely grounded in the video beyond what the masked text alone would suggest.
@@ 5.1 Datasets
LongVideoBench [10] features long instructional clips with QA pairs for extended reasoning, while SUTD-TrafficQA [11] contains short traffic clips with multiple-choice questions on causal and temporal understanding. p-Participants answer one question per each LongVideoBench clip; four questions per each SUTD-TrafficQA clip.
@@ 5.2 User Study Design
Participants are assigned to one of four conditions. In the Video Only condition, they watch the original video without text. In the remaining conditions, they view only a VLM-generated summary: a randomly selected VLM summary (Naive), the top-ranked summary from $k$ response candidates222k=5k=5 in this study. by utility score (Max-U) or by grounding score (Max-G).
@@ Appendix C LearningPaper24 Dataset Curation
OpenReview’s primary areas are often too specific or overlapping, so we remap them into 12 broader and semantically coherent categories. For example, “self-supervised learning” and “representation learning for vision” become Representation Learning, while “probabilistic methods” and “causal inference” are grouped as Probabilistic & Causal Methods. The final taxonomy is shown in Figure˜7. The final dataset includes the following required fields: OpenReview ID, SlidesLive talk ID and URL, TL;DR, abstract, and primary area. We opensource the dataset on Hugging Face.  [^3] https://huggingface.co/datasets/vivianchen98/LearningPaper24
@@ Appendix E User Study Instruction and Interfaces
All participants provide informed consent before participation and are compensated at a rate consistent with Prolific guidelines and institutional standards. The user interface is presented in fig.˜8. Below are the instructions provided to participants for each dataset. Text in parentheses indicates variations between video and text conditions.

### v3

# Shenghui Chen,    Po-han Li11footnotemark: 1,    Sandeep Chinchali,    Ufuk Topcu

@@ Abstract
Many decision-making tasks, where both accuracy and efficiency matter, still require human supervision. For example, tasks like traffic officers reviewing hour-long dashcam footage or researchers screening conference videos can benefit from concise summaries that reduce cognitive load and save time. Yet current vision-language models (VLMs) often produce verbose, redundant outputs that hinder task performance. Existing video caption evaluation depends on costly human annotations and overlooks the summaries’ utility in downstream tasks. We address these gaps with Video-to-text Information Bottleneck Evaluation (VIBE), an annotation-free method that scores VLM outputs using two metrics: grounding (how well the summary aligns with visual content) and utility (how informative it is for the task). VIBE selects from randomly sampled VLM outputs by ranking them according to the two scores to support effective human decision-making. Human studies on LearningPaper24, SUTD-TrafficQA, and LongVideoBench show that summaries selected by VIBE consistently improve performance—boosting task accuracy by up to $61.23\%$ and reducing response time by $75.77\%$ compared to naive VLM summaries or raw video.  [^1] Project Website, Code, and LearningPaper24 Dataset.
@@ 4 Video-to-Text Information Bottleneck Evaluation (VIBE)
We condition both the numerator and denominator with $T_{\mathrm{masked}}$, a masked version of the text, to approximate $\mathbf{I}^{P}(V;T)$. The approximation holds if $T_{\mathrm{masked}}$ is independent of both the video $V$ and the original text $T$. Masking key content makes $T_{\mathrm{masked}}$ effectively uninformative and thus independent. We put the detailed derivation in Appendix I. Following [20], we use tf-idf [37] to identify keywords and replace them with a "<MASK>" token. This masking helps ensure that eq.˜3 reflects the true dependency between $V$ and $T$, rather than exploiting shortcuts from the text itself. Thus, the grounding score measures how much the TL;DR remains genuinely grounded in the video beyond what the masked text alone would suggest.
@@ 5.1 Datasets
LongVideoBench [10] features long instructional clips with QA pairs for extended reasoning, while SUTD-TrafficQA [11] contains short traffic clips with multiple-choice questions on causal and temporal understanding. Participants answer one question per each LongVideoBench clip; four questions per each SUTD-TrafficQA clip.
@@ 5.2 User Study Design
Participants are assigned to one of four conditions. In the Video Only condition, they watch the original video without text. In the remaining conditions, they view only a VLM-generated summary: a randomly selected VLM summary (Naive), the top-ranked summary from $k$ response candidates222k=5k=5 in this study, and the $k$ responses are generated with various temperatures. by utility score (Max-U) or by grounding score (Max-G).
@@ 6 Conclusion
Acknowledgement
This work was supported in part by the National Science Foundation grants No. CNS-1836900, 2148186, the Office of Naval Research (ONR) under Grant No. N00014-22-1-2254, N00014-24-1-2097, the Defense Advanced Research Projects Agency (DARPA) contract FA8750-23-C-1018, and DARPA ANSR: RTXCW2231110. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix C LearningPaper24 Dataset Curation
OpenReview’s primary areas are often too specific or overlapping, so we remap them into 12 broader and semantically coherent categories. For example, “self-supervised learning” and “representation learning for vision” become Representation Learning, while “probabilistic methods” and “causal inference” are grouped as Probabilistic & Causal Methods. The final taxonomy is shown in Figure˜7. The final dataset includes the following required fields: OpenReview ID, SlidesLive talk ID and URL, TL;DR, abstract, and primary area.
@@ Appendix E User Study Instruction and Interfaces
All participants provide informed consent before participation and are compensated at a rate consistent with Prolific guidelines and institutional standards. The user interface is presented in Figure˜8. Below are the instructions provided to participants for each dataset. Text in parentheses indicates variations between video and text conditions.
@@ Appendix G Additional User Study Plots
Appendix H Computation Resource
All experiments were conducted on four NVIDIA RTX 6000 Ada GPUs (48GB VRAM) using the vLLM backend. The system was equipped with an Intel(R) Xeon(R) Gold 6346 CPU @ 3.10GHz, featuring 64 cores (x86_64, 64-bit). This setup handled all models—Qwen-2.5-72B-AWQ, InternVL-2.5-8B-MPO, and InternVL-3-38B—efficiently for both generation and evaluation.
Appendix I Derivation of Mutual Information Approximation
We derive the case where the approximation holds under the stated assumption, starting from the right-hand side of Eq. 3 and Eq. 4. Since both approximations follow nearly identical steps, we present the derivation of grounding score as an example:
$\log\frac{\mathbf{P}(T\mid V,T_{\mathrm{masked}})}{\mathbf{P}(T\mid T_{\mathrm{masked}})}.$
Since $T_{\mathrm{masked}}$ is a masked sentence with all keywords removed, we assume it contains no information and is independent from any other random variables. Thus, we can rewrite the expression as:
$\log\frac{\mathbf{P}(T\mid V,T_{\mathrm{masked}})}{\mathbf{P}(T\mid T_{\mathrm{masked}})}$ $=\log\frac{\mathbf{P}(T,V,T_{\mathrm{masked}})/\mathbf{P}(V,T_{\mathrm{masked}})}{\mathbf{P}(T,T_{\mathrm{masked}})/\mathbf{P}(T_{\mathrm{masked}})}$
$=\log\frac{\mathbf{P}(T,V)\cdot\mathbf{P}(T_{\mathrm{masked}})/(\mathbf{P}(V)\cdot\mathbf{P}(T_{\mathrm{masked}}))}{\mathbf{P}(T)\cdot\mathbf{P}(T_{\mathrm{masked}})/\mathbf{P}(T_{\mathrm{masked}})}$
$=\log\frac{\mathbf{P}(T,V)/\mathbf{P}(V)}{\mathbf{P}(T)}$
$=\log\frac{\mathbf{P}(T\mid V)}{\mathbf{P}(T)}.$
All steps follow from Bayes’ theorem and the assumption of independence.